OpenStudy (anonymous):

Write an equation of the line perpendicular to the line 5x-2y=-3 and passing through the point (2,-1)

OpenStudy (anonymous):

It's an open question

OpenStudy (anonymous):

first you would want to write 5x-2y=-3 in the form of y=mx+c this will help you to find the m (sope) and we know a perpendicular line to another will have - inverse fraction for their slope. then re-write a new y=mx+c with the new perpendicular slope then substitute x as 2 and y as -1

OpenStudy (anonymous):

mx+c or mx+b it is still the same because we now c or b is just a number which is the y intercept

OpenStudy (anonymous):

I got 2x-5y-1=0

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

rewrite 5x-2y=-3 in the form y=mx+b

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

2y=5x+3

OpenStudy (anonymous):

now divide everything by 2 because we're making y the subject of the formula

OpenStudy (anonymous):

5x+3=8

OpenStudy (anonymous):

8/2=4

OpenStudy (anonymous):

2y=4

OpenStudy (anonymous):

2/2 crosses out; 4/2=2

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

2y=5x+3 dividing everything by two will give y= (5/2)x + 3/2

OpenStudy (anonymous):

so, divide the 5x and 3 by 2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

its not finished yet